# High-speed Imaging through Turbulence with Event-based Light Fields

### 14 Mar 2026 v1

@@ 3.1 Modeling Turbulent Event Light Fields
This forward model is bilinear in $S$ and $|h|^{2}$: scene motion (changes in $S$) and turbulence (changes in $h$) are multiplicatively coupled, making their individual contributions ambiguous in the observed intensity $I$. Thus, recovering both from a single observation constitutes an ill-posed blind inverse problem.
Multiple views mitigate this ambiguity. If $N$ observations share a scene $S$ through separate turbulence realizations $\{h^{(i)}\}$, the system accumulates cross-view constraints on the common $S$:
@@
We obtain $N$ views with a light field event camera, with the goal of high speed imaging through turbulence. Following the standard event camera model [e2vid], the intensity of each view is passed through a non-linear logarithmic threshold: an event $(\mathbf{x},t,p)\in\mathcal{E}^{(i)}$ is emitted whenever the absolute log-intensity change $\Delta L^{(i)}(\mathbf{x},t)=|\log I^{(i)}(\mathbf{x},t)-\log I^{(i)}(\mathbf{x},t{-}\Delta t)|$ exceeds a threshold $C$, with polarity $p_{i}=\operatorname{sgn}(\Delta L^{(i)})$.
@@ 3.2 Turbulence Mitigation Model
While the $N$ different views provide critical cross-view constraints on the shared scene $S$, the system remains underdetermined without physical regularization on $h$. Rather than designing explicit priors for this non-convex joint estimation problem, we adopt a data-driven approach: a neural network is trained to directly approximate the inverse mapping from the $N$ event streams to the clean video $V$. By training on a large dataset of simulated turbulence [turbulence_p2s], the network implicitly learns the spatial statistics of anisoplanatic turbulence and the manifold of natural scenes, leveraging cross-view consensus to resolve the bilinear ambiguity.
@@ 3.2.1 Training Details
E2VID is traditionally trained with the calibrated perceptual loss LPIPS instead of mean squared error (MSE) because MSE losses often produce blurry images [e2vid]. The network is trained from scratch for $200$ epochs with a batch size of $8$ using the Adam optimizer [Kingma2015Adam] with a constant learning rate of $10^{-4}$ on a single NVIDIA RTX A6000 GPU.
@@ 4.0.1 Simulated Dataset
A synthetic light field event dataset is constructed by combining a physics-based turbulence simulator (P2S) [turbulence_p2s] with an event simulator (V2E) [v2e], illustrated by Figure 4. Clean intensity frames are sourced from the REDS dataset [Nah_2019_CVPR_Workshops_REDS], which was recorded at 120 fps. The central region is cropped and resized to $256\times 256$, serving as ground truth. For each clip, $N=9$ independent turbulence realizations are generated by running the turbulence simulator with the same parameters but different random seeds for the P2S, producing nine sequences of degraded frames. Each degraded sequence is then passed through the event simulator to produce the corresponding sub-aperture event stream $\mathcal{E}^{(i)}$. The contrast threshold $C$ is uniformly sampled from $[0.1,0.7]$ and the refractory period is set to the default $5$ ms. The temporal window $\Delta T_{k}$ in the voxel grid is set to one frame interval.

### 26 Mar 2026 v2

@@ 3.1 Modeling Turbulent Event Light Fields
This forward model is bilinear in $S$ and $|h|^{2}$: scene motion (changes in $S$) and turbulence (changes in $h$) are multiplicatively coupled, making their individual contributions ambiguous in the observed intensity II. Thus, recovering both from a single observation constitutes an ill-posed blind inverse problem.
Multiple views mitigate this ambiguity. If NN observations share a scene $S$ through separate turbulence realizations $\{h^{(i)}\}$, the system accumulates cross-view constraints on the common $S$:
@@
We obtain NN views with a light field event camera, with the goal of high speed imaging through turbulence. Following the standard event camera model [e2vid], the intensity of each view is passed through a non-linear logarithmic threshold: an event $(\mathbf{x},t,p)\in\mathcal{E}^{(i)}$ is emitted whenever the absolute log-intensity change $\Delta L^{(i)}(\mathbf{x},t)=|\log I^{(i)}(\mathbf{x},t)-\log I^{(i)}(\mathbf{x},t{-}\Delta t)|$ exceeds a threshold $C$, with polarity $p_{i}=\operatorname{sgn}(\Delta L^{(i)})$.
@@ 3.2 Turbulence Mitigation Model
While the NN different views provide critical cross-view constraints on the shared scene $S$, the system remains underdetermined without physical regularization on $h$. Rather than designing explicit priors for this non-convex joint estimation problem, we adopt a data-driven approach: a neural network is trained to directly approximate the inverse mapping from the NN event streams to the clean video $V$. By training on a large dataset of simulated turbulence [turbulence_p2s], the network implicitly learns the spatial statistics of anisoplanatic turbulence and the manifold of natural scenes, leveraging cross-view consensus to resolve the bilinear ambiguity.
@@ 3.2.1 Training Details
E2VID is traditionally trained with the calibrated perceptual loss LPIPS instead of mean squared error (MSE) because MSE losses often produce blurry images [e2vid]. The network is trained from scratch for $200$ epochs with a batch size of 88 using the Adam optimizer [Kingma2015Adam] with a constant learning rate of $10^{-4}$ on a single NVIDIA RTX A6000 GPU.
@@ 4.0.1 Simulated Dataset
A synthetic light field event dataset is constructed by combining a physics-based turbulence simulator (P2S) [turbulence_p2s] with an event simulator (V2E) [v2e], illustrated by Figure 4. Clean intensity frames are sourced from the REDS dataset [Nah_2019_CVPR_Workshops_REDS], which was recorded at 120 fps. The central region is cropped and resized to $256\times 256$, serving as ground truth. For each clip, $N=9$ independent turbulence realizations are generated by running the turbulence simulator with the same parameters but different random seeds for the P2S, producing nine sequences of degraded frames. Each degraded sequence is then passed through the event simulator to produce the corresponding sub-aperture event stream $\mathcal{E}^{(i)}$. The contrast threshold $C$ is uniformly sampled from $[0.1,0.7]$ and the refractory period is set to the default 55 ms. The temporal window $\Delta T_{k}$ in the voxel grid is set to one frame interval.